### v3

@@
Abstract This work tackles a class of optimization problems where fixing some well-chosen combinations of the variables makes the problem substantially easier to solve. We consider that the variables space may be partitioned into subsets that fix these combinations to given values, so the restriction of the problem to any of the partition sets admits a tractable solution. Then, we exhibit a reformulation of the original problem that consists in searching for the partition set index that minimizes the objective value of the solution to the restricted problem. We name partitioned optimization framework (POf) the formalization of this class of problems and this reformulated problem. As we prove in this work, the POf allows solving the original problem by focusing on the reformulated problem: all solutions to the reformulated problem are partition indices for which the solution to the associated restricted problem is also a solution to the original problem. Second, we introduce a derivative-free partitioned optimization method (DFPOm) to efficiently solve problems that fit in the POf. We prove that the reformulated problem is nicely handled by a class of derivative-free optimization (DFO) algorithms named algorithms with a covering step. Then the DFPOm consists in solving the reformulated problem using such DFO algorithm with a covering step to obtain an optimal partition index, and to return the solution to the associated restricted problem as a solution to the initial problem. Finally, we illustrate the DFPOm on a class of problems called composite greybox problems, and we highlight the gain in numerical performance provided by the DFPOm on large-dimensional instances by comparing it to two popular DFO solvers.
@@ 1 Introduction
$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{i}}$)
where $\mathbb{Y}$ is a (possibly infinite-dimensional) Banach space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}$ (possibly discontinuous) has a nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$, and $\Omega\subseteq\mathbb{Y}$ (possibly not closed) intersects $Y$. We consider that $\mathbb{Y}$ may be partitioned so that restricting Problem ($\mathbf{P}_{\texttt{i}}$) to any partition set makes it significantly simpler to solve. We introduce a partitioned optimization framework (POf) to study this context, and derivative-free partition optimization method (DFPOm) to efficiently solve Problem ($\mathbf{P}_{\texttt{i}}$) when it fits in the POf. Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 summarizes our contribution and outlines this work.
@@ 1.1 Illustrative example with a composite greybox problem
Figure 1: The function $\varphi$ from Sections 1.2 and 4.1, minimized efficiently via the POf. For readability, at each $y_{1}\in[-4,4]$ the two graphs actually plot only the values of the quadratic $y_{2}\mapsto\varphi(y_{1},y_{2})$ for $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$.
Minimizing $\varphi$ over $\mathbb{Y}$ is difficult. Yet, $\varphi$ has a simpler expression when $y_{1}$ is fixed: for all $y_{1}\in\mathbb{R}$, the function $y_{2}\mapsto\varphi(y_{1},y_{2})$ is quadratic. We exploit this property by partitioning $\mathbb{Y}$ as $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$, where for all partition set index $x\in\mathbb{X}\triangleq\mathbb{R}$, the partition set $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ is such that the restriction of $\varphi$ to $\mathbb{Y}(x)$ (denoted by $\varphi_{|\mathbb{Y}(x)}$) is simple to minimize. We define an oracle function $\gamma:\mathbb{X}\to\mathbb{Y}$ that, for all $x\in\mathbb{X}$, computes the minimizer of $\varphi_{|\mathbb{Y}(x)}$. Its associated objective function value is denoted by $\Phi(x)\triangleq\varphi(\gamma(x))$. As we highlight in this work, the optimal partition set index $x^{*}$ minimizing $\Phi$ provides a solution to the initial problem as the point $\gamma(x^{*})$. In this example, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is $\gamma(x)\triangleq(x,\sigma(x))$, and its objective function value equals $\Phi(x)=\varepsilon(x)$. The minimizer $x^{*}\triangleq 0$ of $\Phi$ yields $\gamma(x^{*})\triangleq(0,0)$, which is indeed the global solution to the problem.
The POf formalizes the class of optimization problems where a similar approach is possible. We also introduce our DFPOm to compute the optimal partition set $x^{*}$, since the problem of minimizing $\Phi$ is not trivial in general. For completeness, let us also introduce another function the POf relies on, the index function $\chi$, defined for all $y\in\mathbb{Y}$ as the unique index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. In the example above, $\chi(y)\triangleq y_{1}$ for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
@@ 1.2 Motivation for this work
This work introduces the POf as a new optimization framework that leverages inherent structure of some optimization problems to describe their potential solutions, and the DFPOm to exploit the POf in practice. Several applications may enjoy the DFPOm.
A first broad field of applications is optimal control problems [32] with discontinuous Mayer cost. An example consists in optimizing the flight of a parachute so that it lands on a target made of concentric rings with different costs, while minimizing overall efforts. The Mayer cost of the flight is the cost associated to the ring the parachute ends in, which is therefore discontinuous. However, if we impose the landing point of the parachute, the Mayer cost becomes fixed, so the problem becomes usual, and then solvable (at least numerically with standard methods). Thus, we may partition the variables space (the full trajectory of the parachute plus the control to ensure it) according to the landing point. This context was our initial motivation in developing the POf and to consider $\mathbb{Y}$ infinite-dimensional. This study is detailed in our technical report [11] and in the PhD thesis [10, Chapter 7] (in French).
Another rich field is composite greybox problem [21, 23] where, as in Section 1.1, the inputs of the blackbox components of the problem are low-dimensional. Such problems are discussed in Section 4.
@@
Others possible applications may likely connect to the POf, such as almost separable problems [29] and problems that become smooth when a constraint linking the variables is added.
@@ 1.3 Contribution and outline of this work
Our contribution is threefold. First, we state the POf, and we show that we may recover a solution to Problem ($\mathbf{P}_{\texttt{i}}$) from an optimal partition set index. Second, we propose the DFPOm to exploit the POf, using a class of derivative-free optimization (DFO) algorithms to compute an optimal partition set index. Third, we empirically assert the performance of the DFPOm through numerical experiments comparing it against two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{i}}$) directly.
This work is organized as follows. A literature review is provided in Section 2. Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it derives the DFPOm to solve Problem ($\mathbf{P}_{\texttt{i}}$) in Theorem 2. Our third goal follows in Sections 4 and 5. Section 4 shows how the DFPOm solves some composite optimization problems, and that numerical observations match our theoretical claims. Section 5 considers more challenging problem instances, and shows that solving these problems directly with two DFO solvers from the literature is less efficient than our strategy. Finally, Section 6 discusses our work and its possible extensions.
@@ 2 Review of related literature
The POf relates to the frameworks of parametric optimization [8, 30] and bilevel optimization [15]. Indeed, parametric optimization studies problems where the objective function and the feasible set depend on a tuneable parameter. In our case, the parameterized problem is the minimization of $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem is the partition set index $x\in\mathbb{X}$. The local continuity of the parametric solution in discontinuous contexts is studied in [18], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours. Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{i}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem is left to the user (as the indices of the partition set result from the chosen partition of $\mathbb{Y}$). Similarly, bilevel optimization describes problems in which some of the variables are fixed as a solution to a nested subproblem. However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{i}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure. Some references study bilevel nonsmooth problems [25] and apply DFO algorithms [16] but, to the best of our knowledge, none provides results similar to ours.
@@
An important aspect of the POf is its exploitation through the DFPOm. The problem of computing an optimal partition set index is akin to a usual blackbox optimization [1] problem. To solve it, we use DFO algorithms with a covering step [3], since to our best knowledge they are the only algorithms guaranteed to return a local solution in discontinuous contexts under mild assumptions.
@@
Although the DFPOm leverages DFO, we remark that the POf itself targets problems that differ from usual use-cases from DFO. Indeed, the dimension of $\mathbb{Y}$ may be high in Problem ($\mathbf{P}_{\texttt{i}}$), while DFO algorithms usually handle up to a few dozens of variables [6, Section 1.4]. Moreover, the POf requires some information on the problem to define a relevant partition of $\mathbb{Y}$ and ensure the tractability of the minimizer of $\varphi$ over each partition set, which is not available in the blackbox optimization setting that is popular in DFO. The DFO contexts that match the most with the POf are greybox optimization [2] and large-dimension concerns. Greybox optimization considers problems where some information on the problem is available, as in our example in Section 1.2. Regarding large-dimensional DFO, the POf has complementarity with the literature. Indeed, usual techniques rely on dimension-selection [17] that assume active subspaces [14, 34], or random directions [13, 19, 26]. These techniques likely outperform the POf when the objective function has a small sensitivity with respect to most of the variables but a high sensitivity with respect to the few others. Yet, they likely fail on cases where the objective function mostly depends on a few complex combinations of the variables instead of all variables directly. In contrast, this last context is favourable to the POf. As illustrated in Section 5, the DFPOm outperforms usual DFO algorithms on four problems with this property.
@@ 3 Formal framework and related theoretical results
This section considers Problem ($\mathbf{P}_{\texttt{i}}$) in its general form (that is, $\varphi$ is not limited to its expression from the introductory example). Section 3.1 formalizes the POf and its associated properties, and Section 3.2 details the DFPOm that we use to exploit the POf and highlights the resulting convergence analysis. The proofs of our claims are left to Section 3.3.
Notation: We denote by $\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$, $\overline{\mathbb{N}}\triangleq\mathbb{N}\cup\{+\infty\}$, $\mathbb{R}^{*}\triangleq\mathbb{R}\setminus\{0\}$ and $\overline{\mathbb{N}}^{*}\triangleq\overline{\mathbb{N}}\setminus\{0\}$. Consider two normed spaces $\mathbb{A}$ and $\mathbb{B}$. A set $A\subseteq\mathbb{A}$ is said to be ample [3] if $A\subseteq\mathrm{cl}(\mathrm{int}(A))$ (where $\mathrm{cl}$ and $\mathrm{int}$ denote respectively the closure and interior operators), and to be a continuity set of a function $f:\mathbb{A}\to\overline{\mathbb{R}}$ if $f(a)$ is finite for all $a\in A$ and $f_{|A}$ is continuous. For all collections $(A_{i})_{i=1}^{N}$ of subsets of $\mathbb{A}$, with $N\in\overline{\mathbb{N}}^{*}$, their union is denoted by $\sqcup_{i=1}^{N}A_{i}$ when the sets are pairwise disjoint. For all sequences $(b^{k})_{k\in\mathbb{N}}\in\mathbb{B}^{\mathbb{N}}$, we denote by $\mathrm{acc}((b^{k})_{k\in\mathbb{N}})$ the set of all accumulation points of $(b^{k})_{k\in\mathbb{N}}$. For all functions $g:\mathbb{A}\to\mathbb{B}$, all sets $\mathcal{A}\subseteq\mathbb{A}$ and all points $a\in\mathbb{A}$, we denote by $g(\mathcal{A})\triangleq\{g(a):a\in\mathcal{A}\}$ and by $\mathcal{ACC}(g;a)$ the union of $\mathrm{acc}((g(a^{k}))_{k\in\mathbb{N}})$ over all sequences $(a^{k})_{k\in\mathbb{N}}\in\mathbb{A}^{\mathbb{N}}$ converging to $a$. For all functions $f:\mathbb{A}\to\overline{\mathbb{R}}$ and all sets $A\subseteq\mathbb{A}$, we denote by $\overline{f_{A}}:\mathbb{A}\to\overline{\mathbb{R}}$ the extreme barrier function [4] mapping all $a\in\mathbb{A}$ to $f(a)$ if $a\in A$ and to $\infty$ if $a\notin A$. In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{i}}$).

### v5

@@
Abstract This work tackles a class of optimization problems in which fixing some well-chosen combinations of the variables makes the problem substantially easier to solve. We consider that the variables space may be partitioned into subsets that fix these combinations to given values, so that the restriction of the problem to any of the partition sets admits a tractable solution. Then, we exhibit a reformulation of the problem that consists in searching for the partition set index that minimizes the objective value of the solution to the restricted problem. We name partitioned optimization framework (POf) the formalization of this class of problems and this reformulation process. As we prove in this work, the POf allows solving the original problem by focusing on the reformulated problem: all solutions to the reformulated problem are partition indices for which the solution to the associated restricted problem is also a solution to the original problem. Second, we introduce a derivative-free partitioned optimization method (DFPOm) to efficiently solve problems that fit in the POf. We prove that the reformulated problem is nicely handled by derivative-free optimization (DFO) algorithms with a covering step. Then the DFPOm consists in solving the reformulated problem using such DFO algorithm with a covering step to obtain an optimal partition index, and to return the solution to the associated restricted problem as a solution to the initial problem. Finally, we illustrate how the POf allows solving some classes of problems. We first focus on an infinite-dimensional case, by solving analytically an optimal control problem that challenges standard methods from the literature. Then, we apply the DFPOm on a class of finite-dimensional problems called composite greybox problems, and we highlight the gain in numerical performance provided by the DFPOm by comparing it to two popular DFO solvers.
@@ 1 Introduction
$\underset{\begin{array}[]{c}y\in\mathbb{Y}\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{ini}}$)
where $\mathbb{Y}$ is a (possibly infinite-dimensional) Banach space, and $\varphi:\mathbb{Y}\to\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$ (possibly discontinuous) has a nonempty effective domain $Y\triangleq\{y\in\mathbb{Y}:\varphi(y)\neq\pm\infty\}$ intersecting $\Omega\subseteq\mathbb{Y}$ (which is possibly not closed). We consider that $\mathbb{Y}$ may be partitioned so that restricting Problem ($\mathbf{P}_{\texttt{ini}}$) to any partition set makes it significantly simpler to solve. We introduce a partitioned optimization framework (POf) to study this context, and a derivative-free partition optimization method (DFPOm) to efficiently solve Problem ($\mathbf{P}_{\texttt{ini}}$) when it fits in the POf. Section 1.1 illustrates the POf, Section 1.2 discusses our motivation to develop it, and Section 1.3 outlines this work and our contribution.
@@ 1.1 Illustrative example with a composite greybox problem
Figure 1: The function $\varphi$ from Sections 1.2 and 5.1, minimized efficiently via the POf. For readability, at each $y_{1}\in[-4,4]$ the two graphs actually plot only the values of the quadratic $y_{2}\mapsto\varphi(y_{1},y_{2})$ for $y_{2}\in[\sigma(y_{1})-1.3,\ \sigma(y_{1})+1.3]$.
Minimizing $\varphi$ over $\mathbb{Y}$ is difficult. Yet, $\varphi$ has a simpler expression when $y_{1}$ is fixed: for all $y_{1}\in\mathbb{R}$, the function $y_{2}\mapsto\varphi(y_{1},y_{2})$ is quadratic. We exploit this property by partitioning $\mathbb{Y}$ as $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$, where for all partition set index $x\in\mathbb{X}\triangleq\mathbb{R}$, the partition set $\mathbb{Y}(x)\triangleq\{x\}\times\mathbb{R}$ is such that the restriction of $\varphi$ to $\mathbb{Y}(x)$ (denoted by $\varphi_{|\mathbb{Y}(x)}$) is simple to minimize. We define an oracle function $\gamma:\mathbb{X}\to\mathbb{Y}$ that, for all $x\in\mathbb{X}$, computes the minimizer of $\varphi_{|\mathbb{Y}(x)}$. Its associated objective function value is denoted by $\Phi(x)\triangleq\varphi(\gamma(x))$. As we highlight in this work, the optimal partition set index $x^{*}$ minimizing $\Phi$ provides a solution to the initial problem as the point $\gamma(x^{*})$. In this example, for all $x\in\mathbb{X}$, the minimizer of $\varphi_{|\mathbb{Y}(x)}$ is $\gamma(x)\triangleq(x,\sigma(x))$, and its objective function value equals $\Phi(x)=\varepsilon(x)$. The minimizer $x^{*}\triangleq 0$ of $\Phi$ yields $\gamma(x^{*})=(0,0)$, which is indeed the global solution to the problem.
The POf formalizes the class of optimization problems where a similar approach is possible. We also introduce our DFPOm to compute the optimal partition set index $x^{*}$, since the problem of minimizing $\Phi$ is not trivial in general. For completeness, let us also introduce another function the POf relies on, the index function $\chi$, defined for all $y\in\mathbb{Y}$ as the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. In the example above, $\chi(y)\triangleq y_{1}$ for all $y\triangleq(y_{1},y_{2})\in\mathbb{Y}$, since $y\in\mathbb{Y}(x)$ only when $x=y_{1}$.
@@ 1.2 Motivation for this work
This work introduces the POf as a new optimization framework that leverages inherent structure of some optimization problems to describe their potential solutions, and the DFPOm to exploit the POf in practice. Several applications may enjoy the POf and be solved efficiently through the DFPOm.
A first broad field of applications is optimal control problems [36] with discontinuous Mayer cost. An example consists in optimizing the flight of a parachute so that it lands on a target made of concentric rings with different costs, while minimizing overall efforts. The Mayer cost of the flight is the cost associated to the ring the parachute ends in, which is therefore discontinuous. However, if we impose the landing point of the parachute, the Mayer cost becomes fixed, so the problem becomes smooth, and then solvable (at least numerically with standard methods). Thus, we may partition the variables space (the full trajectory of the parachute plus the control to ensure it) according to the landing point. This context was our initial motivation in developing the POf and to consider $\mathbb{Y}$ infinite-dimensional. In Section 4 an example of discontinuous optimal control problem is provided. It challenges standard methods but is analytically solved using the POf. A broader application of the POf in discontinuous contexts of optimal control theory is given in the PhD thesis [11, Chapter 7] (in French), and in our technical report [10].
Another rich field is those of composite greybox problems [24, 23] where, as in Section 1.1, the inputs of the blackbox components of the problem are low-dimensional. Several examples of such problems are discussed in Section 5.
@@
Other possible applications may likely connect to the POf, such as almost separable problems [33] and problems that become smooth when a constraint linking the variables is added.
@@ 1.3 Contribution and outline of this work
Our contribution is fourfold. First, we state the POf, and we show that we may recover a solution to Problem ($\mathbf{P}_{\texttt{ini}}$) from an optimal partition set index. Second, we propose the DFPOm to exploit the POf, using a class of derivative-free optimization (DFO) algorithms to compute an optimal partition set index. Third, we substantiate our infinite-dimensional considerations by solving with the POf a discontinuous optimal control problem that challenges usual techniques from the literature. Fourth, we empirically assert the performance of the DFPOm through numerical experiments ensuring that our theoretical analysis is reflected in practice, and comparing the DFPOm against two state-of-the-art DFO solvers solving Problem ($\mathbf{P}_{\texttt{ini}}$) directly.
This work is organized as follows. A literature review is provided in Section 2. Then, Section 3 addresses our first two goals: it formalizes the POf and highlights its properties in Theorem 1, and it derives the DFPOm to solve Problem ($\mathbf{P}_{\texttt{ini}}$) in Theorem 2. Our third goal follows in Section 4, which uses Theorem 1 to solve analytically a discontinuous version of a classical optimal control problem on which standard methods cannot be applied directly. Finally, our fourth goal is addressed in Sections 5 and 6. Section 5 shows how the DFPOm solves some finite-dimensional composite optimization problems, and that numerical observations match our theoretical claims from Theorem 2. Section 6 considers more challenging problem instances (in higher dimension), and shows that solving these problems directly with two DFO solvers from the literature is less efficient than our strategy. Finally, Section 7 discusses our work and its possible extensions.
@@ 2 Review of related literature
The POf relates to the frameworks of parametric optimization [34] and bilevel optimization [16]. Indeed, parametric optimization studies problems where the objective function and the feasible set depend on a tuneable parameter. In our case, the parameterized problem is to minimize $\varphi$ over $\mathbb{Y}(x)\cap\Omega$, where the parameter of the problem is the partition set index $x\in\mathbb{X}$. The local continuity of the parametric solution in discontinuous contexts is studied in [19], but, to the best of our knowledge, no reference in parametric optimization claims results similar to ours. Moreover, the terminology partitioned optimization highlights that Problem ($\mathbf{P}_{\texttt{ini}}$) is non-parametric and that the nature of the so-called parameter in the restricted problem is left to the user (as the partition set indices result from the chosen partition of $\mathbb{Y}$). Similarly, bilevel optimization describes problems in which some of the variables are fixed as a solution to a nested subproblem. However, the difference in terminology highlights a contextual difference since Problem ($\mathbf{P}_{\texttt{ini}}$) has only one level and, unlike in usual bilevel problems, there is no conflictual min-max structure. Some references study bilevel nonsmooth problems [26] and apply DFO algorithms [17] but, to the best of our knowledge, none provides results similar to ours.
@@
An important aspect of the POf is its exploitation through the DFPOm. The problem of computing an optimal partition set index is akin to a usual blackbox optimization [1] problem. To solve it, we use DFO algorithms with a covering step [3] since, to our best knowledge, they are the only algorithms with guarantees to return a local solution in discontinuous contexts under mild assumptions.
@@
Although the DFPOm leverages DFO, we remark that the POf itself targets problems that differ from usual use-cases from DFO. Indeed, the dimension of $\mathbb{Y}$ may be high in Problem ($\mathbf{P}_{\texttt{ini}}$), while DFO algorithms usually handle up to a few dozens of variables [6, Section 1.4]. Moreover, the POf requires some information on the problem to define a relevant partition of $\mathbb{Y}$ and ensure the tractability of the minimizer of $\varphi$ over each partition set, which is not available in the blackbox optimization setting that is popular in DFO. The DFO contexts that match the most with the POf are greybox optimization [2] and large-dimension concerns. Greybox optimization considers problems where partial information on the problem is available, as in our example in Section 1.1. Regarding large-dimensional DFO, the POf has complementarity with the literature. Indeed, usual techniques rely on dimension-selection [18] that assume active subspaces [15, 38], or random directions [13, 20, 29]. These techniques likely outperform the POf when the objective function has a small sensitivity with respect to most of the variables but a high sensitivity with respect to the few others. Yet, they likely fail on cases where the objective function mostly depends on a few complex combinations of the variables instead of all variables directly. In contrast, this last context is favourable to the POf. As illustrated in Section 6, the DFPOm outperforms usual DFO algorithms on four problems with this property.
@@ 3 Formal framework and related theoretical results
This section considers Problem ($\mathbf{P}_{\texttt{ini}}$) in its general form (that is, $\varphi$ is not limited to its expression from Section 1.1). Section 3.1 formalizes the POf and its associated properties, and Section 3.2 details the DFPOm that we use to exploit the POf and highlights the resulting convergence analysis. The proofs of our claims are left to Section 3.3.
Notation: We denote by $\overline{\mathbb{R}}\triangleq\mathbb{R}\cup\{\pm\infty\}$, $\overline{\mathbb{N}}\triangleq\mathbb{N}\cup\{+\infty\}$, $\mathbb{R}^{*}\triangleq\mathbb{R}\setminus\{0\}$ and $\overline{\mathbb{N}}^{*}\triangleq\overline{\mathbb{N}}\setminus\{0\}$. Consider two normed spaces $\mathbb{A}$ and $\mathbb{B}$. A set $A\subseteq\mathbb{A}$ is said to be ample [3] if $A\subseteq\mathrm{cl}(\mathrm{int}(A))$ (where $\mathrm{cl}$ and $\mathrm{int}$ denote respectively the closure and interior operators), and to be a continuity set of a function $f:\mathbb{A}\to\overline{\mathbb{R}}$ if $f(a)$ is finite for all $a\in A$ and $f_{|A}$ is continuous. For all collections $(A_{i})_{i=1}^{N}$ of subsets of $\mathbb{A}$, with $N\in\overline{\mathbb{N}}^{*}$, their union is denoted by $\sqcup_{i=1}^{N}A_{i}$ when the sets are pairwise disjoint. For all sequences $(b^{k})_{k\in\mathbb{N}}\in\mathbb{B}^{\mathbb{N}}$, we denote by $\mathrm{acc}((b^{k})_{k\in\mathbb{N}})$ the set of all accumulation points of $(b^{k})_{k\in\mathbb{N}}$. For all functions $g:\mathbb{A}\to\mathbb{B}$, all sets $\mathcal{A}\subseteq\mathbb{A}$ and all points $a\in\mathbb{A}$, we denote by $g(\mathcal{A})\triangleq\{g(a):a\in\mathcal{A}\}$ and by $\mathcal{ACC}(g;a)$ the union of $\mathrm{acc}((g(a^{k}))_{k\in\mathbb{N}})$ over all sequences $(a^{k})_{k\in\mathbb{N}}\in\mathbb{A}^{\mathbb{N}}$ converging to $a$. For all functions $f:\mathbb{A}\to\overline{\mathbb{R}}$ and all sets $A\subseteq\mathbb{A}$, we denote by $\overline{f_{A}}:\mathbb{A}\to\overline{\mathbb{R}}$ the extreme barrier function [4] mapping all $a\in\mathbb{A}$ to $f(a)$ if $a\in A$ and to $\infty$ if $a\notin A$. In particular, $\overline{\varphi_{\Omega}}$ denotes the extreme barrier function associated to $\varphi$ and $\Omega$ from Problem ($\mathbf{P}_{\texttt{ini}}$).